# From Sufficiency to Reflection: Reinforcement-Guided Thinking Quality in Retrieval-Augmented Reasoning for LLMs

### v1

@@ 4.3 Reward Design
Thinking reward measures the quality of the thinking part of the trajectory. The locally deployed LLM is prompted to evaluate logical soundness, alignment with retrieved content, error awareness, and concise accuracy. It outputs a score in [0,1][0,1][ 0 , 1 ]:
@@ 4.4 Optimization Strategy
Difficulty-aware Resampling. Despite employing a mixed reward mechanism, we observed that different rollout groups can yield very different raw rewards but almost identical normalized advantages (e.g., [0.8,0.85,0.9,0.95,1.0][0.8,0.85,0.9,0.95,1.0][ 0.8 , 0.85 , 0.9 , 0.95 , 1.0 ] vs. [0.0,0.05,0.1,0.15,0.2][0.0,0.05,0.1,0.15,0.2][ 0.0 , 0.05 , 0.1 , 0.15 , 0.2 ]). This phenomenon causes simple problems to receive the same optimization emphasis as hard problems. Following Zhang and Zuo (2025), we introduce a difficulty-aware advantage reweighting strategy. For each sample, we estimate the problem difficulty by computing the average sufficient-reward score across all its rollouts, denoted as $R_{S}^{\text{avg}}$. The weight function is defined as
@@ Appendix B Experimental Setup for Preliminary Studies
We define three cases: (1) Overthinking: the model produces too many reasoning steps, meaning that the gold answer could already be inferred at some step $R_{i}$ with $i<n$. (2) Good thinking: the model obtains sufficient content exactly before the final reasoning step, i.e., the gold answer can only be inferred at step $R_{n+1}$. (3) Underthinking: the model stops reasoning without obtaining sufficient content in $RD$, and still outputs an answer. Since Joren et al. (2025) has shown enabling LLMs to judge whether the provided context is adequate for answering a question is effective, we directly input $Q$, $RD$, and the gold answer into an LLM to assess whether $Q$ and $RD$ together are sufficient to derive the correct answer (see prompt in App.J). Importantly, we include the gold answer in the input, unlike what Joren et al. (2025) does. which uses only $Q$ and $RD$. This is because we found that omitting the answer leads to more errors: when $RD$ is actually insufficient, the model is more likely to incorrectly judge it as sufficient.
@@ Appendix C Implementation Details and Baselines
[…]
@@ 4. Clarity and Precision – Is it concise, relevant, and non-redundant?
Question: Who is Gaius Rubellius Blandus’s father-in-law? Golden Answer: Drusus Julius Caesar Naive GRPO: <think> To find out who Gaius Rubellius Blandus’s father-in-law is, I need to first determine who his wife was. I’ll search for information about Gaius Rubellius Blandus to find details about his family, particularly his wife. </think>
@@
Question: Where was the place of burial of John De Ferrers, 1st Baron Ferrers of Chartley’s father? Golden Answer: Staffordshire Naive GRPO: <think> To find out where John de Ferrers, 1st Baron Ferrers of Chartley’s father was buried, I need to first determine who his father was. I’ll search for information about John de Ferrers, 1st Baron Ferrers of Chartley to find his father’s name. </think>
@@
Question: Who was born first, Dennis E. Nolan or Humberto Anguiano? Golden Answer: Dennis E. Nolan Naive GRPO: <think> To determine who was born first between Dennis E. Nolan and Humberto Anguiano, I need to find out their birth dates. I’ll first search for Dennis E. Nolan’s birth date. </think>
@@
From the search results, I found that Dennis E. Nolan was a United States Army general. Now, I’ll search for Humberto Anguiano’s birth date. </think>

### v2

@@ 4.3 Reward Design
Thinking reward measures the quality of the thinking part of the trajectory. The locally deployed LLM is prompted to evaluate logical soundness, alignment with retrieved content, error awareness, and concise accuracy. It outputs a score in $[0,1]$:
@@ 4.4 Optimization Strategy
Difficulty-aware Resampling. Despite employing a mixed reward mechanism, we observed that different rollout groups can yield very different raw rewards but almost identical normalized advantages (e.g., $[0.8,0.85,0.9,0.95,1.0]$ vs. $[0.0,0.05,0.1,0.15,0.2]$). This phenomenon causes simple problems to receive the same optimization emphasis as hard problems. Following Zhang and Zuo (2025), we introduce a difficulty-aware advantage reweighting strategy. For each sample, we estimate the problem difficulty by computing the average sufficient-reward score across all its rollouts, denoted as $R_{S}^{\text{avg}}$. The weight function is defined as
@@ Appendix B Experimental Setup for Preliminary Studies
We define three cases: (1) Overthinking: the model produces too many reasoning steps, meaning that the gold answer could already be inferred at some step $R_{i}$ with $i<n$. (2) Good thinking: the model obtains sufficient content exactly before the final reasoning step, i.e., the gold answer can only be inferred at step $R_{n+1}$. (3) Underthinking: the model stops reasoning without obtaining sufficient content in $RD$, and still outputs an answer. Since Joren et al. (2025) has shown enabling LLMs to judge whether the provided context is adequate for answering a question is effective, we directly input $Q$, RDRD, and the gold answer into an LLM to assess whether $Q$ and $RD$ together are sufficient to derive the correct answer (see prompt in App.J). Importantly, we include the gold answer in the input, unlike what Joren et al. (2025) does. which uses only $Q$ and $RD$. This is because we found that omitting the answer leads to more errors: when $RD$ is actually insufficient, the model is more likely to incorrectly judge it as sufficient.
@@ Appendix C Implementation Details and Baselines
For all baselines and TIRESRAG-R1, we use Qwen-2.5-3B (base and instruct variants) as the backbone. Following R1-Searcher, we use the 2019 English Wikipedia as the external knowledge source, and BGE-large-en-v1.5 as the retrieval engine. Our training framework is built upon OpenRLHF Hu et al. (2025b), and we use FlashRAG Jin et al. (2025b) for evaluation. The total batch size is 108, with a learning rate of $2\times 10^{-6}$. We generate 5 rollout samples per input for reward estimation and set temperature = 1 during rollout to encourage exploration. The KL coefficient $\beta$ is set to 0, the number of iterations per batch $\mu$ = 2, and the clipping parameter $\epsilon=0.2$. Top-$k$ during retrieval is set to 5. In the difficulty-aware resampling strategy, the hyperparameters are set to $A=0.4$, $B=1.5$, $\rho_{0}=0.75$, and $k=10.0$. For the consistency penalty, we set $\lambda_{p}=0.1$. In the overall reward computation, the weights are configured as $w_{t}=0.6$, $w_{s}=0.3$, and $w_{r}=0.3$. All models are trained on 3 NVIDIA H200 GPUs: 2 GPUs are allocated for policy optimization, and 1 GPU is dedicated to rollout inference via vLLM Kwon et al. (2023).
@@ 4. Clarity and Precision – Is it concise, relevant, and non-redundant?
Question: Who is Gaius Rubellius Blandus’s father-in-law? Golden Answer: Drusus Julius Caesar Naive GRPO: <think> To find out who Gaius Rubellius Blandus’s father-in-law is, I need to first determine who his wife was. I’$l$ search for information about Gaius Rubellius Blandus to find details about his family, particularly his wife. </think>
@@
Question: Where was the place of burial of John De Ferrers, 1st Baron Ferrers of Chartley’s father? Golden Answer: Staffordshire Naive GRPO: <think> To find out where John de Ferrers, 1st Baron Ferrers of Chartley’s father was buried, I need to first determine who his father was. I’$l$ search for information about John de Ferrers, 1st Baron Ferrers of Chartley to find his father’s name. </think>
@@
Question: Who was born first, Dennis E. Nolan or Humberto Anguiano? Golden Answer: Dennis E. Nolan Naive GRPO: <think> To determine who was born first between Dennis E. Nolan and Humberto Anguiano, I need to find out their birth dates. I’$l$ first search for Dennis E. Nolan’s birth date. </think>
@@
From the search results, I found that Dennis E. Nolan was a United States Army general. Now, I’$l$ search for Humberto Anguiano’s birth date. </think>